OpenStudy (anonymous):

Solve the equation below for x: 2(x-1)=5(x+1)-3 A. x = -4/3 B. x = 0 C. x = -3/4 D. x = 5/2

OpenStudy (anonymous):

you can be begin by distributiing : 2(x-1)=5(x+1)-3 you can be begin by distributiing : 2x-2=5x+5-3 Now just collect like terms